OpenStudy (anonymous):

Find the width of a rectangle with a perimeter of 90 and a length of 15. (1 point)A90 B15 C45 D30

OpenStudy (anonymous):

2l+2w=p plug in the information you have. 2(15)=2w=90 solve for w.

OpenStudy (anonymous):

p=2*length+2*width 90=2*15+2*x x=(90-30)/2=30 Therefore it should be D

OpenStudy (anonymous):

2(15)+2(w)=90 sorry!